OpenStudy (anonymous):

can someone help me please.. I need the equation solution for :y=9+.08(x) and y=15+.03(x)

OpenStudy (anonymous):

set then equal to each other, write \[9+.08x=15+.03x\] then perhaps \[900+8x=1500+3x\] \[5x=600\] \[x=120\] substitute to find y